OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x - sin x

OpenStudy (mathstudent55):

Use a double angle identity.

OpenStudy (anonymous):

Could you please elaborate @mathstudent55

OpenStudy (zpupster):

cos(2x) = cos^2(x) - sin^2(x) = 2 cos^2(x) - 1 = 1 - 2 sin^2(x)

OpenStudy (anonymous):

These are the answers cos2x - sin2x - sin x cos2x - sin3x cos2x + sin2x + sin x cos2x - 3 sin x

OpenStudy (anonymous):

cos2x - sin2x - sin x should be the answer if that 2 means square?

OpenStudy (anonymous):

It does. Why do you see that the answer is A?

OpenStudy (anonymous):

\[\cos 2x= \cos^2x - \sin^2 x\]

OpenStudy (anonymous):

Ok i'll put that as my answer. Thanks again!